# Greedy Algorithm for Structured Bandits: A Sharp Characterization of Asymptotic Success / Failure*

### v1

@@ 5 Interactive decision-making with arbitrary feedback
Now let’s analyze the rounds in which the decoy policy $\pi_{\mathtt{dec}}$ is chosen. Let $t(j)$ be the $j$-th such round, $j\in\mathbb{N}$. Informally, we’d like to argue that throughout all these rounds, the likelihood of the decoy model $M_{\mathtt{dec}}$ grows faster than that of any other model $M\in\mathcal{M}$. To this end, we consider $\Psi_{j}(M):=\Delta\ell_{t(j)}(M_{\mathtt{dec}})-\Delta\ell_{t(j)}(M)$ and the event
@@
Here, we restrict to $j>N_{0}/2$ to ensure that events $E_{1},E_{2}$ concern disjoint sets of events, and therefore can be treated as independent. We observe that $E_{1}\cap E_{2}$ implies that in each round $t>T_{0}$, $\mathcal{L}_{t}(M_{\mathtt{dec}})>\mathcal{L}_{t}(M)$ for any model $M\in\mathcal{M}_{\mathtt{other}}$, and so $\mathtt{GreedyMLE}$ chooses the decoy policy.
@@
Finally, we argue that $E_{2}$ happens with positive probability. W.l.o.g., we can assume that the outcomes $(r_{t},o_{t})$ in all rounds $t=t(j)$, $j\in\mathbb{N}$ are drawn in advance from an “outcome tape”. [^16] Namely: we have an “outcome tape” whose entries are drawn independently from $M^{*}(\pi_{\mathtt{dec}})$, and $(r_{t(j)},o_{t(j)})$ is defined as the $j$-entry of this tape, for all $j\in\mathbb{N}$. We leverage Eq. (5.5) once again for this argument. Indeed, $\Psi_{j}(M)=0$ for every model $M\in\mathcal{M}$ that fully coincides with $M_{\mathtt{dec}}$ on the decoy policy $\pi_{\mathtt{dec}}$, so we only need to worry about the models $M\in\mathcal{M}$ for which this is not the case. Then $\sum_{i\in[j]}\operatornamewithlimits{\mathbb{E}}\left[\,\Psi_{i}(M)\,\right]\geq j\cdot\Gamma$. We obtain $\sum_{i\in[j]}\operatornamewithlimits{\mathbb{E}}\left[\,\Psi_{i}(M)\,\right]\geq 0$ with positive-constant probability by appropriately applying concentration separately for each $j>N_{0}/2$ and taking a union bound.
@@ 6 Examples
Let us instantiate our characterization for several well-studied reward structures from bandits literature. We consider linear, Lipschitz, and (one-dimensional) polynomial structures, for bandits as well as contextual bandits. All structures are discretized to ensure finiteness, as required in our framework. The discretization is consistent across different reward functions, in the sense that all functions take values in the same (discrete) set $\mathcal{R}$, with $|\mathcal{R}|\ll|\mathcal{F}|$. This prevents a trivial form of self-identifiability that could arise if each reward function $f$ were discretized independently and inconsistently, resulting in some $f(a)$ values being unique and making $f$ self-identifiable solely due to the discretization strategy specific to $f$ [^17] For example, consider an instance $(f^{*},\mathcal{F})$ being not self-identifiable, with its decoy $f_{\mathtt{dec}}\in\mathcal{F}$ satisfying $f^{*}(a_{\mathtt{dec}})=f_{\mathtt{dec}}(a_{\mathtt{dec}})=0.5$. Now, suppose we discretize $f^{*}(a_{\mathtt{dec}})$ using discretization step 0.1 and discretize $f_{\mathtt{dec}}(a_{\mathtt{dec}})$ using discretization step 0.2. After this modification, $f^{*}(a_{\mathtt{dec}})$ and $f_{\mathtt{dec}}(a_{\mathtt{dec}})$ would no longer be equal, and self-identifiability could occur.
@@
Proof 6.2.
Let $j\in[d]$ be a coordinate with the smallest $|\theta^{*}_{j}|$. Choose arm $a_{\mathtt{dec}}\in\left\{\,-1,1\,\right\}^{d}$ with $(a_{\mathtt{dec}})_{i}=\mathrm{sign}(\theta^{*}_{i})$ for all coordinates $i\neq j$, and flipping the sign for $i=j$, $(a_{\mathtt{dec}})_{j}=-\mathrm{sign}(\theta^{*}_{j})$. Note that $\langle\theta^{*},a_{\mathtt{dec}}\rangle=\|\theta^{*}\|_{1}-2\,\min_{i\in[d]}|\theta^{*}_{i}|\in[d\varepsilon,d]$.
Now, for any given $\alpha\in[d\varepsilon,\,d]_{\varepsilon}$ and any sign vector $v\in\{-1,1\}^{d}$, there is $\theta\in\Theta$ such that $\|\theta\|_{1}=\alpha$ and its signs are aligned as $\mathrm{sign}(\theta)=v$. Thus, there exists $\theta_{\mathtt{dec}}\in\Theta$ such that $\|\theta_{\mathtt{dec}}\|_{1}=\langle\theta^{*},a_{\mathtt{dec}}\rangle$ and $\mathrm{sign}(\theta_{\mathtt{dec}})=a_{\mathtt{dec}}$. Note that $a_{\mathtt{dec}}$ is the best arm for $\theta_{\mathtt{dec}}$. Moreover, $\langle\theta_{\mathtt{dec}},a_{\mathtt{dec}}\rangle=\|\theta_{\mathtt{dec}}\|_{1}=\langle\theta^{*},a_{\mathtt{dec}}\rangle<\|\theta^{*}\|_{1}=\langle\theta^{*},a^{*}\rangle$. So, $\theta_{\mathtt{dec}}$ is a decoy for $\theta^{*}$.
@@ 8 Conclusions
We study the greedy algorithm in structured bandits and characterize its asymptotic success vs failure in terms of a simple partial-identifiability property of the problem structure. Our characterization holds for arbitrary finite structures and extends to bandits with contexts and/or auxiliary feedback. In particular, we find that the greedy algorithm succeeds only if the problem is intrinsically “easy” for any algorithm which satisfies a mild non-degeneracy condition.
@@
We identify three directions for further work. First, extend our characterization to infinite function/model classes. Then one can no longer rely on function-gap being strictly positive, which necessitates a more subtle analysis. Second, extend our characterization to approximate greedy algorithms, stemming either from approximate regression or from human behaviorial biases. A class of approximately-greedy algorithms, representing myopic human behavior under behavioral biases, was studied in Banihashem et al. (2023a, b), but only for unstructured multi-armed bandits. Third, while our “asymptotic” perspective enables a sharp characterization over reward structures, stronger guarantees are desirable for a particular reward structure. More concretely: negative results with failure probabilities guaranteed to be not extremely small, and/or positive results with regret rates that are guaranteed to not have extremely large “constant factors”. Such guarantees are only known for a few specific reward structures. [^23] Specifically, prior work provides strong positive guarantees for linear contextual bandits with smoothed/diverse contexts (Kannan et al., 2018; Bastani et al., 2021; Raghavan et al., 2023), and strong negative guarantees for unstructured multi-armed bandits and Bayesian bandits with priors independent across arms (Banihashem et al., 2023a, b).
@@ Appendix A $\mathtt{StructuredCB}$ with tie-breaking
Proof of Theorem A.2. The proof of Theorem 4.3(b) mostly carries over, with the following minor modifications. Let $\mathcal{A}^{*}_{\mathtt{dec}}(x)=\operatornamewithlimits{argmax}_{a\in\mathcal{A}}f_{\mathtt{dec}}(x)$ be the set of optimal arms for the decoy $f_{\mathtt{dec}}$ for a given context $x$. The two events $E_{1}$ and $E_{2}$ (as originally defined eq. 4.3 and eq. 4.4) will be modified to be invoked on all decoy context-arm pairs.
@@
Analyzing the probability for event $E_{1}$ still follows from Lemma F.3. Analyzing the probability for event $E_{2}$ follows from Lemma F.5, but with the choice of $\sigma$ will be chosen as $\sigma=\Theta(\Gamma(f_{\mathtt{dec}})/\sqrt{\ln(|\mathcal{X}|K)})$, and we still have $\Pr[E_{2}]\geq 0.9$.
Appendix B $\mathtt{StructuredMAB}$ Success: Proof of Theorem 3.3(a)
@@
Lemma B.3.
Assume the event in Lemma B.1, then we have the upper bound on $\mathtt{MSE}_{\tau}$ for each round $\tau\in[t]$.
@@
Lemma B.5.
Assume the event in Lemma B.1. The number of times any suboptimal arm is chosen cannot exceed $T^{\prime}$ rounds, where $T^{\prime}$ is some parameter with $T^{\prime}=(K/\Gamma(f^{*})^{2})\cdot O(\log t)$.
Proof B.6.
@@
Appendix C $\mathtt{StructuredMAB}$ Failure: Proof of Theorem 3.3(b)
@@
For any $x$ in the interval $[f_{\mathtt{dec}}(a)-\Gamma(f_{\mathtt{dec}})/2,f_{\mathtt{dec}}(a)-\Gamma(f_{\mathtt{dec}})/2]$, by boundedness of mean reward, we have
@@
Then, the density of $x$ at any point on the interval $[f_{\mathtt{dec}}(a)-\Gamma(f_{\mathtt{dec}})/2,f_{\mathtt{dec}}(a)-\Gamma(f_{\mathtt{dec}})/2]$ is at least
@@
For some appropriately chosen $\sigma=\Theta(\Delta)$, we have event $E_{2}$ happens with probability at least
@@
Theorem 3.3(b) directly from the above lemmas.
Appendix D $\mathtt{StructuredMAB}$ Failure: deterministic decoy arm
We obtain an improved failure probability, as per Remark 3.6, when the decoy arm $a_{\mathtt{dec}}$ gives deterministic rewards (and all other arms give unit-variance Gaussian rewards, as before).  [^24] This is a departure from $\mathtt{StructuredCB}$ as defined in Section 2, where all arms have unit-Gaussian rewards. Note that $\mathtt{Greedy}$ does not know that $a_{\mathtt{dec}}$ is deterministic – as it does not input any knowledge about reward noise. The failure probability depends on the ”standard” instance-gap between the best and the second-best arm:
$\Delta(f)=\max_{a}f(a)-\max_{a^{\prime}:f(a^{\prime})<f(a)}f(a^{\prime}).$
We obtain essentially the same guarantee as in Theorem 3.3(b), with function-gap $\Gamma(f_{\mathtt{dec}})$ replaced with instance-gap $\Delta(f_{\mathtt{dec}})$.
Theorem D.1.
Suppose the warm-up data consists of one sample for each arm. Assume $f^{*}$ has a decoy $f_{\mathtt{dec}}\in\mathcal{F}$, with decoy arm $a_{\mathtt{dec}}$. Suppose $a_{\mathtt{dec}}$ produces deterministic rewards, whereas all other arms produce unit-variance Gaussian rewards. Then with some probability $p_{\mathtt{dec}}>0$ it holds that $\mathtt{Greedy}$ chooses $a_{\mathtt{dec}}$ for all rounds $t\in(T_{0},\infty)$. We can lower-bound $p_{\mathtt{dec}}$ by $\Omega((\Delta(f_{\mathtt{dec}})/\sqrt{K}))^{K-1}$.
The first part of the theorem stating that $\mathtt{Greedy}$ only chooses $a_{\mathtt{dec}}$ follows from the below two lemmas. The second part lower bounding the value of $p_{\mathtt{dec}}$ is by directly examining the gaussian density.
We focus on the following event.
$E_{4}:=\left\{\forall a\neq a_{\mathtt{dec}},\absolutevalue{\bar{r}_{\tt warm}(a)-f_{\mathtt{dec}}(a)}<\Delta(f_{\mathtt{dec}})/\sqrt{16K}\right\}.$
Lemma D.2.
For any round $t$, if up until round $t$ only $a_{\mathtt{dec}}$ is chosen, under event $E_{4}$ we have $\mathtt{MSE}(f_{\mathtt{dec}})<\Delta(f_{\mathtt{dec}})^{2}/16$.
Proof D.3.
For square loss from arms $a\neq a_{\mathtt{dec}}$, we have
$\sum_{a\neq a_{\mathtt{dec}}}N_{t}(a)(\bar{r}_{t}(a)-f_{\mathtt{dec}}(a))<(K-1)(\Delta(f_{\mathtt{dec}})/\sqrt{16K})^{2}<\Delta(f_{\mathtt{dec}})^{2}/16.$
For the arm $a_{\mathtt{dec}}$, we have with constant probability, the following holds:
$N_{t}(a_{\mathtt{dec}})(\bar{r}_{t}(a)-f_{\mathtt{dec}}(a))=0.$
Lemma D.4.
For any round $t$, if up until round $t$ only decoy arm $a_{\mathtt{dec}}$ is chosen, then in the next round the decoy arm $a_{\mathtt{dec}}$ is chosen.
Proof D.5.
Consider any function $f$ with $f(a)>f(a_{\mathtt{dec}})$ for some arm $a$. We argue that $f$ cannot possibly be the $\mathtt{MSE}$ minimizer. Since we know that under event $E_{4}$
|$\absolutevalue{\bar{r}_{\tt warm}(a)-f_{\mathtt{dec}}(a)}<\Delta(f_{\mathtt{dec}})/\sqrt{16K}<\Delta(f_{\mathtt{dec}})/4,$
together with $f(a)>f(a_{\mathtt{dec}})$, we know that |$\absolutevalue{\bar{r}_{\tt warm}(a)-f(a)}>\Gamma(f_{\mathtt{dec}})/2$. Then
$\mathtt{MSE}_{t}(f)>(\bar{r}_{1}(a)-f(a))^{2}>\Delta^{2}(f_{\mathtt{dec}})/4.$
Therefore, any such $f$ cannot be the $\mathtt{MSE}$ minimizer this round. In other words, the $\mathtt{MSE}$ minimizer must have $f_{t}(a_{\mathtt{dec}})>f(a)$, and $a_{\mathtt{dec}}$ will continue to be chosen.
Appendix E $\mathtt{StructuredCB}$ Success: Proof of Theorem 4.3(a)

### v2

@@ 5 Interactive decision-making with arbitrary feedback
Let’s analyze the rounds in which the decoy policy $\pi_{\mathtt{dec}}$ is chosen. Let $t(j)$ be the $j$-th such round, $j\in\mathbb{N}$. We’d like to argue that throughout all these rounds, the likelihood of the decoy model $M_{\mathtt{dec}}$ grows faster than that of any other model $M\in\mathcal{M}$. To this end, consider event
$E_{2}:=\{\;\forall j>N_{0}/2,\;\forall M\in\mathcal{M}_{\mathtt{other}},\quad{\textstyle\sum_{i\in[j]}}\Psi_{i}(M)\geq 0\;\},$
@@
where $\Psi_{j}(M):=\Delta\ell_{t(j)}(M_{\mathtt{dec}})-\Delta\ell_{t(j)}(M)$. Here, we restrict to $j>N_{0}/2$ to ensure that $E_{1},E_{2}$ concern disjoint sets of events, and hence are independent. $E_{1}\cap E_{2}$ implies that in each round $t>T_{0}$, $\mathcal{L}_{t}(M_{\mathtt{dec}})>\mathcal{L}_{t}(M)$ for any model $M\in\mathcal{M}_{\mathtt{other}}$, and so $\mathtt{GreedyMLE}$ chooses the decoy policy.
@@
Finally, we argue that $E_{2}$ happens with positive probability. W.l.o.g., the outcomes $(r_{t},o_{t})$ in all rounds $t=t(j)$, $j\in\mathbb{N}$ are drawn in advance from an “outcome tape”. [^14] Its entries $j\in\mathbb{N}$ are drawn independently from $M^{*}(\pi_{\mathtt{dec}})$, and $(r_{t(j)},o_{t(j)})$ is defined as the $j$-entry. We leverage Eq. (5.5) once again. Indeed, $\Psi_{j}(M)=0$ for every model $M\in\mathcal{M}$ that fully coincides with $M_{\mathtt{dec}}$ on the decoy policy $\pi_{\mathtt{dec}}$, so we only need to worry about the models $M\in\mathcal{M}$ for which this is not the case. Then $\sum_{i\in[j]}\operatornamewithlimits{\mathbb{E}}\left[\,\Psi_{i}(M)\,\right]\geq j\cdot\Gamma$. We obtain $\sum_{i\in[j]}\operatornamewithlimits{\mathbb{E}}\left[\,\Psi_{i}(M)\,\right]\geq 0$ with positive-constant probability by appropriately applying concentration separately for each $j>N_{0}/2$ and taking a union bound.
6 Structured bandits with an Infinite Function Class
We obtain a partial characterization for $\mathtt{StructuredMAB}$, which handles an arbitrary infinite function class $\mathcal{F}$ and yields better constants compared to Theorem 3.3. The success of $\mathtt{Greedy}$ requires a stronger notion of self-identifiability: approximately fixing the expected reward of a suboptimal arm identifies it as suboptimal. The failure of $\mathtt{Greedy}$ requires a stronger notion of a decoy function, which must lie in the “interior” of $\mathcal{F}$. The characterization is “partial” in the sense that it does not yield a full dichotomy. However, the boundary between success and failure is controlled by a tunable “margin” parameter $\varepsilon>0$, which can be made arbitrarily small (and optimized based on the instance).
Definition 6.1.
A problem instance $(f^{*},\mathcal{F})$ is $\varepsilon$-self-identifiable, $\varepsilon\geq 0$, if any suboptimal arm $a$ of $f^{*}$ is suboptimal for any reward function $f\in\mathcal{F}$ with $|f(a)-f^{*}(a)|\leq\varepsilon$. An $\varepsilon$-interior of $\mathcal{F}$, $\mathtt{int}(\mathcal{F},\varepsilon)$ is the set of all functions $f\in\mathcal{F}$, such that any reward function $f^{\prime}$ with $\|f^{\prime}-f\|_{2}\leq\varepsilon$ is also in $\mathcal{F}$.
For a “continuous” function class such as linear functions or Lipschitz functions, $\mathtt{int}(\mathcal{F},\varepsilon)$ typically includes all but an $O(\varepsilon)$-fraction of $\mathcal{F}$. The choice of the $\ell_{2}$ norm in the definition of $\varepsilon$-interior is not essential: any $\ell_{p}$ norm suffices. We provide the main theorem below; see proof in Appendix F.
Theorem 6.2.
Fix a problem instance $(f^{*},\mathcal{F})$ of $\mathtt{StructuredMAB}$ with an infinite function class $\mathcal{F}$ (but a finite action set $\mathcal{A}$). For any $\varepsilon>0$ (which can be optimized based on $f^{*}$):
  • (a)
    If the problem instance is $\varepsilon$-self-identifiable, then $\mathtt{Greedy}$ (with any warm-up data) satisfies $\operatornamewithlimits{\mathbb{E}}\left[\,R(t)\,\right]\leq T_{0}+(K/\varepsilon)^{2}\cdot O(\log t)$ for each round $t\in\mathbb{N}$.
  • (b)
    Suppose the warm-up data consists of one sample for each arm. Assume $f^{*}$ has a decoy $f_{\mathtt{dec}}\in\mathtt{int}(\mathcal{F},\varepsilon)$, with decoy arm $a_{\mathtt{dec}}$. Then with some probability $p_{\mathtt{dec}}>0$ it holds that $\mathtt{Greedy}$ chooses $a_{\mathtt{dec}}$ for all rounds $t\in(T_{0},\infty)$. We can lower-bound $p_{\mathtt{dec}}$ by $e^{-O(K^{2}/\varepsilon^{2})}$.
This result mirrors Theorem 3.3, with the function gap replaced by $\varepsilon$, allowing for instance-dependent optimization of $\varepsilon$ and tighter bounds. The proof for part (a) carries over with simple modifications. In contrast, proving part (b) is considerably more subtle. In the infinite case, $\mathtt{Greedy}$ may not get stuck on a single reward function—it could almost surely switch among infinitely many. The key insight is that such fluctuations need not impact the arm choice: even as the predictor changes, the greedy selection may remain fixed. The proof exploits this decoupling, constructing events where the algorithm persistently selects a decoy arm, even as the greedy predictors continue to evolve.
@@ 7 Examples
Let us instantiate our characterization for several well-studied reward structures from bandits literature. We consider linear, Lipschitz, and (one-dimensional) polynomial structures, for bandits as well as contextual bandits. All reward structures in this section are discretized to ensure finiteness, as required for our complete characterization in Sections 3, 4 and 5. (While our partial characterization in Section 6 handles infinite reward structures, a secondary goal of this section is to illustrate how common infinite reward structures can be meaningfully discretized so that the complete finite-structure results become directly applicable.) The discretization is consistent across different reward functions, in the sense that all functions take values in the same (discrete) set $\mathcal{R}$, with $|\mathcal{R}|\ll|\mathcal{F}|$. This prevents a trivial form of self-identifiability that could arise if each reward function $f$ were discretized independently and inconsistently, resulting in some $f(a)$ values being unique and making $f$ self-identifiable solely due to the discretization strategy specific to $f$ [^15] For example, consider an instance $(f^{*},\mathcal{F})$ being not self-identifiable, with its decoy $f_{\mathtt{dec}}\in\mathcal{F}$ satisfying $f^{*}(a_{\mathtt{dec}})=f_{\mathtt{dec}}(a_{\mathtt{dec}})=0.5$. Now, suppose we discretize $f^{*}(a_{\mathtt{dec}})$ using discretization step 0.1 and discretize $f_{\mathtt{dec}}(a_{\mathtt{dec}})$ using discretization step 0.2. After this modification, $f^{*}(a_{\mathtt{dec}})$ and $f_{\mathtt{dec}}(a_{\mathtt{dec}})$ would no longer be equal, and self-identifiability could occur.
@@
Proof 7.2.
Let $j\in[d]$ be a coordinate with the smallest $|\theta^{*}_{j}|$. Choose arm $a_{\mathtt{dec}}\in\left\{\,-1,1\,\right\}^{d}$ with $(a_{\mathtt{dec}})_{i}=\mathrm{sign}(\theta^{*}_{i})$ for all coordinates $i\neq j$, and flipping the sign for $i=j$, $(a_{\mathtt{dec}})_{j}=-\mathrm{sign}(\theta^{*}_{j})$. Note that $\innerproduct{\theta^{*}}{a_{\mathtt{dec}}}=\|\theta^{*}\|_{1}-2\,\min_{i\in[d]}|\theta^{*}_{i}|\in[d\varepsilon,d]$.
Now, for any given $\alpha\in[d\varepsilon,\,d]_{\varepsilon}$ and any sign vector $v\in\{-1,1\}^{d}$, there is $\theta\in\Theta$ such that $\|\theta\|_{1}=\alpha$ and its signs are aligned as $\mathrm{sign}(\theta)=v$. Thus, there exists $\theta_{\mathtt{dec}}\in\Theta$ such that $\|\theta_{\mathtt{dec}}\|_{1}=\innerproduct{\theta^{*}}{a_{\mathtt{dec}}}$ and $\mathrm{sign}(\theta_{\mathtt{dec}})=a_{\mathtt{dec}}$. Note that $a_{\mathtt{dec}}$ is the best arm for $\theta_{\mathtt{dec}}$. Moreover, $\innerproduct{\theta_{\mathtt{dec}}}{a_{\mathtt{dec}}}=\|\theta_{\mathtt{dec}}\|_{1}=\innerproduct{\theta^{*}}{a_{\mathtt{dec}}}<\|\theta^{*}\|_{1}=\innerproduct{\theta^{*}}{a^{*}}$. So, $\theta_{\mathtt{dec}}$ is a decoy for $\theta^{*}$.
@@ 9 Conclusions
We study the greedy algorithm in structured bandits and characterize its asymptotic success vs failure in terms of a simple partial-identifiability property of the problem structure. Our characterization holds for arbitrary finite structures and extends to bandits with contexts and/or auxiliary feedback. In particular, we find that the greedy algorithm succeeds only if the problem is intrinsically “easy” for any algorithm which satisfies a mild non-degeneracy condition. We also provide a partial characterization for $\mathtt{StructuredMAB}$ with infinite reward structures (and finite action sets).
@@
We identify three directions for further work. First, extend our positive and negative results to infinite action sets (and infinite function/model classes). Ideally one would like to obtain a full characterization, like in the finite case. Second, extend our characterization to approximate greedy algorithms, stemming either from approximate regression or from human behaviorial biases. A class of approximately-greedy algorithms, representing myopic human behavior under behavioral biases, was studied in Banihashem et al. (2023a, b), but only for unstructured multi-armed bandits. Third, while our “asymptotic” perspective enables a sharp characterization over reward structures, stronger guarantees are desirable for a particular reward structure. More concretely: negative results with failure probabilities guaranteed to be not extremely small, and/or positive results with regret rates that are guaranteed to not have extremely large “constant factors”. Such guarantees are only known for a few specific reward structures. [^21] Specifically, prior work provides strong positive guarantees for linear contextual bandits with smoothed/diverse contexts (Kannan et al., 2018; Bastani et al., 2021; Raghavan et al., 2023), and strong negative guarantees for unstructured multi-armed bandits and Bayesian bandits with priors independent across arms (Banihashem et al., 2023a, b).
@@ Appendix A $\mathtt{StructuredCB}$ with tie-breaking
Proof of Theorem A.2. The proof of Theorem 4.3(b) mostly carries over, with the following minor modifications. Let $\mathcal{A}^{*}_{\mathtt{dec}}(x)=\operatornamewithlimits{argmax}_{a\in\mathcal{A}}f_{\mathtt{dec}}(x)$ be the set of optimal arms for the decoy $f_{\mathtt{dec}}$ for a given context $x$. The two events $E_{1}$ and $E_{2}$ (as originally defined eq. D.1 and eq. D.2) will be modified to be invoked on all decoy context-arm pairs.
@@
Analyzing the probability for event $E_{1}$ still follows from Lemma D.11. Analyzing the probability for event $E_{2}$ follows from Lemma D.13, but with the choice of $\sigma$ will be chosen as $\sigma=\Theta(\Gamma(f_{\mathtt{dec}})/\sqrt{\ln(|\mathcal{X}|K)})$, and we still have $\Pr[E_{2}]\geq 0.9$.
Appendix B Novelty of self-identifiability
We argue that self-identifiability is a novel notion. Specifically, we compare it to (i) knowing the optimal value, and (ii) Graves-Lai coefficient being 00.
First, one could ask if self-identifiability is equivalent to knowing the value of the best arm. However, the former does not imply the latter. Consider the simple example =$\mathcal{F}=\left\{(3,1),(2,1)\right\}$. Both functions are self-identifiable in $\mathcal{F}$, but clearly the optimal value differs.
Second, consider the Graves-Lai coefficient (Graves and Lai, 1997; Wagenmaker and Foster, 2023). Let us define it formally, for the sake of completeness. Consider $\mathtt{DMSO}$, as defined in Section 5, with model class $\mathcal{M}$. Let
$\Delta(\pi|M)=f(\pi_{M}|M)-f(\pi|M)$
be the suboptimality gap for model $M$ and policy $\pi$, where $\pi_{M}$ is the optimal policy for $M$. Let $\mathcal{M}^{\mathtt{alt}}$ be the set of models that disagree with $M$ on the optimal policy:
$\mathcal{M}^{\mathtt{alt}}(M):=\left\{M^{\prime}\in\mathcal{M}|\pi_{M}\neq\pi_{M^{\prime}}\right\}.$
Now, the Graves-Lai coefficient is defined as
$\mathsf{GLC}(\mathcal{M},M)=\inf_{\eta\in\mathbb{R}_{+}^{\Pi}}\left\{\sum_{\pi\in\Pi}\eta_{\pi}\Delta(\pi|M)\ |\ \forall M^{\prime}\in\mathcal{M}^{\mathtt{alt}}(M):\sum_{\pi\in\Pi}\eta_{\pi}D_{\rm KL}\left(M(\pi)\|M^{\prime}(\pi)\right)\geq 1\right\}.$
Intuitively, the Graves-Lai coefficient measures the “verification” cost of verifying whether a given function $f^{*}$ (or a given model $M^{*}$ in the $\mathsf{DMSO}$ setting) is indeed the true model. The Graves-Lai coefficient being 0 implies that the learner can ascertain that $f^{*}$ or $M^{*}$ is indeed the true model by simply executing the set of optimal policies $\Pi(f^{*})$ or $\Pi(M^{*})$.
Now, one could ask if self-identifiability is equivalent to $\mathsf{GLC}(\mathcal{M},M)=0$. We observe that this is not the case: the two notions are incomparable. For a counterexample, consider $\mathtt{StructuredMAB}$ with two arms and =$\mathcal{F}=\left\{(2,1),(0.5,1)\right\}$. Problem instance $f^{*}=(0.5,1)$ is self-identifiable, since revealing the sub-optimal arm as having reward 0.5 immediately rules out $(2,1)$ as being the true model. But the $\mathsf{GLC}>0$, since to ascertain $(0.5,1)$ as being the true model one necessarily has to choose the 1st arm and experiment. On the other hand, one can see $f^{*}=(2,1)$ is not self-identifiable but has $\mathsf{GLC}=0$. In this example, $\mathtt{Greedy}$ succeeds when $\mathsf{GLC}>0$ (larger $\mathsf{GLC}$ suggests larger regret of the optimal algorithm in $\mathsf{GLC}$-based theory) but fails when $\mathsf{GLC}=0$ (lower $\mathsf{GLC}$ suggests lower regret of the optimal algorithm in $\mathsf{GLC}$-based theory)! Hence $\mathsf{GLC}$ does not capture the per-instance behavior of $\mathtt{Greedy}$.
However, $\mathsf{GLC}$ has some connection to our machinery. Namely, if $\mathsf{GLC}(\mathcal{F},f)=0$ for some reward function $f$, then $f$ necessarily cannot be a decoy for any other reward function $f^{*}$. That said, $\mathsf{GLC}(\mathcal{F},f)$ provides no information about whether $f$ itself admits a decoy. We believe that $\mathsf{GLC}$ precisely characterizes the asymptotic performance of the optimal algorithm (Graves and Lai, 1997; Wagenmaker and Foster, 2023), whereas self-identifiability precisely captures the asymptotic behavior of $\mathtt{Greedy}$ —a generally suboptimal algorithm.
Appendix C $\mathtt{StructuredMAB}$ characterization: Proof of Theorem 3.3
C.1 $\mathtt{StructuredMAB}$ Success: Proof of Theorem 3.3(a)
@@ Appendix B Novelty of self-identifiability
Lemma C.3.
Assume the event in Lemma C.1, then we have the upper bound on $\mathtt{MSE}_{\tau}$ for each round $\tau\in[t]$.
@@
Lemma C.5.
Assume the event in Lemma C.1. The number of times any suboptimal arm is chosen cannot exceed $T^{\prime}$ rounds, where $T^{\prime}$ is some parameter with $T^{\prime}=(K/\Gamma(f^{*})^{2})\cdot O(\log t)$.
Proof C.6.
@@
C.2 $\mathtt{StructuredMAB}$ Failure: Proof of Theorem 3.3(b)
@@
For any $x$ in the interval $[f_{\mathtt{dec}}(a)-\Gamma(f_{\mathtt{dec}})/2,f_{\mathtt{dec}}(a)+\Gamma(f_{\mathtt{dec}})/2]$, by boundedness of mean reward, we have
@@
Then, the density of $x$ at any point on the interval $[f_{\mathtt{dec}}(a)-\Gamma(f_{\mathtt{dec}})/2,f_{\mathtt{dec}}(a)+\Gamma(f_{\mathtt{dec}})/2]$ is at least
@@
For some appropriately chosen $\sigma=\Theta(\Gamma(f_{\mathtt{dec}}))$, we have event $E_{2}$ happens with probability at least
@@
Theorem 3.3(b) directly follows from the above lemmas.
Appendix D $\mathtt{StructuredCB}$ characterization: Proof of Theorem 4.3
We start with a proof sketch, and proceed with full proofs.
Part (a). Directly applying the proof technique from the MAB case results in a regret bound that is linear in $|\Pi|=K^{X}$. Instead, we apply a potential argument and achieve regret bound that is polynomial in $KX$. First, by a standard concentration inequality, we upper-bound the loss for $f^{*}$ as $\mathtt{MSE}_{t}(f^{*})\leq\widetilde{O}(KX)$ with high probability. Then, we use self-identifiability to argue that if in some round $t$ of the main stage some suboptimal policy $\pi$ is chosen, there must exist some context-arm pair $(x,\pi(x))$ that is “under-explored”: appeared less than $\widetilde{O}(XK/\Gamma^{2}(f^{*})$ times. This step carefully harnesses the structure of the contextual bandit problem. Finally, we introduce a well-designed potential function (see Lemma D.7) that tracks the progress of learning over time. This function increases whenever a suboptimal policy is executed on an under-explored context-action pair, allowing us to bound the total number of times any suboptimal policy is executed. A key challenge is that while the second step guarantees the existence of an “under-explored” context-arm pair, it does not ensure that the context actually appears when the associated policy is chosen. We address this using a supermartingale argument and the fact that each context arrives with probability at least $p_{0}$ in each round. Combining these steps, we upper-bound the expected number of times $\mathtt{Greedy}$ selects a suboptimal policy, and we bound the final expected regret via the regret decomposition lemma.
Part (b). As in the MAB case, we define event $E_{1}$ to ensure that the warm-up data misidentifies $f_{\mathtt{dec}}$ as the true reward function, and event $E_{2}$ that the empirical rewards of the decoy policy are tightly concentrated. The definitions are modified to account for contexts:
$E_{1}$ ={|$=\left\{\,\absolutevalue{\bar{r}_{{\mathtt{warm}}}(x,a)-f^{\dagger}(x,a)}<\Gamma(f^{\dagger})/2\quad\text{for each $x\in\mathcal{X}$ and arm $a\neq\pi^{\dagger}(x)$}\,\right\},$ (D.1)
$E_{2}$ ={|$=\left\{\,\absolutevalue{\bar{r}_{t}(x,\pi^{\dagger}(x))-f^{*}(x,\pi^{\dagger}(x))}<\Gamma(f^{\dagger})/2\quad\text{for each $x\in\mathcal{X}$ and round $t>T_{0}$}\,\right\}.$ (D.2)
A decoy context-arm pair $(x,a)$ is one with $a=\pi_{\mathtt{dec}}(x)$. $E_{1}$ concerns the single warm-up sample for each non-decoy context-arm pair. $E_{2}$ asserts that the empirical rewards are concentrated for all decoy context-arm pairs (and all rounds throughout the main stage). The two events are independent, as they concern non-overlapping sets of context-arm pairs. $\mathtt{Greedy}$ always chooses the decoy arm when $E_{1},E_{2}$ happen. To lower-bound $\Pr\left[\,E_{1}\cap E_{2}\,\right]$, invoke independence, analyze each event separately.
D.1 $\mathtt{StructuredCB}$ Success: Proof of Theorem 4.3(a)